OpenStudy (ohmath123):

determine how many different values can arise by inserting one pair of parenthesis into the given expression 19-12-8-2

OpenStudy (anonymous):

(19-12)-8-2=-17 19-(12-8)-2=-25 19-12-(8-2)=-13 19-(12-8-2)=17 Answer: 4 different values

OpenStudy (anonymous):

(19-12-8)-2 = -3 so 5 different ways in total